F=X2+y2-z=1
Fine the unit vector is normal to this surface at P(1,1,1)
1
Expert's answer
2019-10-31T11:15:59-0400

Given surface is-


"f(x,y,z)=x^2+y^2-z=1"

So,

The normal at point (1,1,1) will be-

Direction ratios of normal will be-


"(\frac{1}{1},\frac{1}{1},\frac{1}{1})"

"Or\\(1,1,1)"

Unit normal vector =


"\frac{1}{\sqrt{3}}\hat{i}+\frac{1}{\sqrt{3}}\hat{j}+\frac{1}{\sqrt{3}}\hat{k}"